(x - 4)/2 - 1/5 = (7x + 1)/20
Found 2 solutions by sdmmadam@yahoo.com, atif.muhammad:
Answer by sdmmadam@yahoo.com(530)

Solve for x: (x - 4)/2 - 1/5 = (7x + 1)/20
(x - 4)/2 - 1/5 = (7x + 1)/20
Multiplying by 20 which is the lcm of 2,5 and 20
10(x-4)-4X(1)=(7x+1)
10x-40-4=7x+1
10x-44=7x+1
10x-7x=1+44 (grouping like terms, here side changed and hence sign changed)
3x=45
x=45/3=15
Answer: x=15
Verification: putting x=15 in the given equation,we get LHS = (55-2)/10 = 53/10
and RHS = (7X15+1)/20 = (105+1)/20 = 106/20 = 53/10

Answer by atif.muhammad(135)